# Preprocessing Disks for Convex Hulls, Revisited

### v1

@@ 2.1 Disk Characterization
First, observe that a guaranteed disk $D\in\mathcal{R}$ must be a boundary disk, as by Observation 2.1, $D$ must lie outside $E$, or equivalently outside the convex hull of $\mathcal{R}\setminus\{D\}$.
Second, a potential disk $D\in\mathcal{R}$ must be unstable, as $D$ being potential means there is both a realization of $\mathcal{R}$ where $D$ is realized as a vertex of the convex hull and a realization of $\mathcal{R}$ where it is not a vertex of the convex hull, and thus the combinatorial structure depends on the realization of $D$ [^3] The definition of stable does not directly prohibit a scenario where for some fixed realization of $\mathcal{R}\setminus\{D\}$, any realization of $D$ is always a vertex of the convex hull, while for some other fixed realization of $\mathcal{R}\setminus\{D\}$, any realization of $D$ is never a vertex of the convex hull. However, such a scenario is impossible due to the continuity of our disk based uncertain regions.
Finally, an impossible disk must be both interior and stable. Clearly it is stable as its realization is never a vertex of the convex hull and thus does not affect its combinatorial structure. It is interior as by Observation 2.1 it lies inside $I$ which lies strictly inside the conex hull of $\mathcal{R}$.
The above observation implies there are only 5 possible types of disks are (see Figure 4).
@@
Proof 2.7 (Proof of Lemma 2.6).
First, the boundary disks can be easily detected in $O(n\log n)$ time by computing the convex hull of the disks [12]. Second, Observation 2.1 allows us to classify the disks into guaranteed, potential, and impossible disks using a simple $O(n\log n)$ algorithm [29]. By Lemma 2.4 all impossible interior disks are stable. It remains to distinguish the stable from the unstable guaranteed boundary disks.
Let $D\in\mathcal{R}$ be a stable guaranteed boundary disk. Let $B(\mathcal{R})=\{B_{1},\ldots,B_{k}\}$ be the ordered set of boundary disks. As $D$ is a boundary disks, $D=B_{i}$ for some $i$, where $B_{i}$ has two adjacent strips, namely $\mathrm{CH}(B_{i-1},B_{i})$ and $\mathrm{CH}(B_{i},B_{i+1})$.
First, observe that no disk in $\mathcal{R}\setminus\{B_{i-1},B_{i}\}$ can properly intersect the strip $\mathrm{CH}(B_{i-1},B_{i})$ (and the same holds for $\mathrm{CH}(B_{i},B_{i+1})$). To see this, suppose there is at least one such disk, and that its realization in inside this strip. Now if $B_{i-1}$ and $B_{i}$ are realized at the points in the disks that are extreme in the direction orthogonal to the spine of this strip and exterior to the hull, then any realization of any other disk intersecting this strip cannot be a vertex of the convex hull. Conversely, if $B_{i-1}$ and $B_{i}$ are realized at the opposite extremes, then at least one of the other disks whose realization is in this strip will be a vertex of the convex hull.
By the same reasoning, $B_{i-1}$ must be a guaranteed disk (and similarly for $B_{i+1}$). Specifically, if $B_{i-1}$ is not guaranteed it intersects the convex hull of $\mathcal{R}\setminus\{B_{i-1}\}$, and because $B_{i-1}$ is a boundary disk, it in fact intersects a strip of $\mathcal{R}\setminus\{B_{i-1}\}$. Moreover, this strip must have $B_{i}$ as one of its defining disks as the strip $\mathrm{CH}(B_{i-1},B_{i})$ had no intersecting disks. Thus by the same reasoning used to argue the strip $\mathrm{CH}(B_{i-1},B_{i})$ had no intersecting disks, $B_{i-1}$ cannot intersect this strip, and so must be guaranteed.
So in summary, in order for a guaranteed boundary disk to be stable, its adjacent strips must be empty and its neighbors on the strip hull must be guaranteed. Furthermore, it is clear that any guaranteed boundary disk with these properties is stable as then the realizations of $B_{i-1},B_{i},B_{i+1}$ all appear on the convex hull as they are all guaranteed, and no vertex can occur between the realization of $B_{i-1}$ and $B_{i}$ (nor $B_{i-1}$ and $B_{i+1}$), effectively blocking $D=B_{i}$ from affecting the combinatorial structure of the rest of the convex hull.
As for the running time for finding the stable guaranteed boundary disk, as discussed above we can already determine all boundary disks, and for each such disks whether it and its neighbors are guaranteed, in $O(n\log n)$ time. Thus we only determine whether the adajacent strips are empty, though this can be done for all stips in $O(n\log n)$ time using Lemma 5.19.
@@ 2.2 Quarter Hulls
Note that the convex hull is cyclic in nature: it has no well-defined start or end point, and in the presence of uncertainty, it may be difficult even to designate an arbitrary disk as start point, as there may not be any guaranteed disks.
To make our lives easier, we opt to focus on only one quarter of the convex hull: the part from the rightmost point on the hull to the topmost point on the hull (i.e. the positive quadrant). Indeed, the idea of splitting the computation of the convex hull into separate pieces is well established [5], and choosing quarters will have the added benefit that within such a quarter, no vertex can have an internal angle smaller than $90^{\circ}$, which will aid our geometric arguments later.
@@
In the remainder of this work, we also adapt the notation established in Section 2.1 to the notion of quarter hulls. That is, a disk will be called guaranteed or impossible when it is guaranteed or impossible to contribute to the quarter hull, and it will be stable when it does not influence the combinatorial structure of the quarter hull.
Notably, observe that the total number of unstable disks for the entire convex hull is at most a constant number more than the sum of the number of unstable disks in the four quarter hull problems; hence, dividing the problem into quarters does not jeopardize our aim to achieve sublinear reconstruction times.
@@ 2.3 Supersequences
Let $\mathcal{R}$ be a set of $n$ regions in the plane. A quarter-hull-supersequence of $\mathcal{R}$ is a sequence $\Xi$ of (possibly reoccurring) disks of $\mathcal{R}$ such that, for any point set $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$, the quarter hull of $P$ is a subsequence of $\Xi$.
We will focus on the case where $\mathcal{R}$ is a set of unit disks of bounded ply $\Delta$ in this work, where the ply of a set of regions is the maximum number of regions that any point in the plane is contained in. Replacing the disks in a quarter-hull-supersequence by their points, we now obtain a sequence of points that contains all points of the quarter hull of $P$ in the correct order. Since it is not clear if the convex hull of any such sequence can be computed in linear time (Open Problem 2.15), we additionally introduce smooth sequences.
For two points $p,q\in\mathbb{R}^{2}$, let $\nwarrow(p,q)$ be the signed distance from $p$ to $q$ projected on a line with slope $-1$; that is,
@@
Essentially, a smooth supersequence may have points that are out of order, but the extent in which they can be out of order is limited by the parametesr $\alpha$ and $\beta$, which makes the reconstruction problem significantly easier.
@@
2.4 Statement of results
We now state our main results for preprocessing unit disks in the plane for the convex hull. Our first three theorems imply that the classical result of $O(n\log n)$ preprocessing for $O(n)$ reconstruction (as in Figure 1) can indeed be replicated for unit disks of bounded ply using supersequences.
@@ 2.4 Statement of results
Let $\mathcal{R}$ be a set of $n$ unit disks in the plane of ply $\Delta$. There exists a $(3\sqrt{2},12\Delta)$-smooth quarter-hull-supersequence $\Xi$ of $\mathcal{R}$ of size $O(\Delta^{2}n)$.
@@
Let $\mathcal{R}$ be a set of $n$ unit disks in the plane of ply $\Delta$. A $(3\sqrt{2},12\Delta)$-smooth quarter-hull-supersequence $\Xi$ of $\mathcal{R}$ of size $O(\Delta^{2}n)$ can be computed in $O(\Delta n(\Delta+\log n))$ time.
@@
Let $\mathcal{R}$ be a set of $n$ unit disks. Given an $(\alpha,\beta)$-smooth quarter-hull-supersequence $\Xi$ of $\mathcal{R}$, and given a point set $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$, the quarter hull of $P$ can be computed in $O(\alpha\beta|\Xi|)$ time.
The next theorem implies that we can also obtain sublinear reconstruction using this method.
Theorem 2.13.
Let $\mathcal{R}$ be a set of $n$ unit disks. We are given an $(\alpha,\beta)$-smooth quarter-hull-supersequence $\Xi$ of $\mathcal{R}$ in which all its stable disks have been marked, and let $\mu$ denote the number of unstable disks in $\Xi$. Then given a point set $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$, the quarter hull of $P$ can be computed in $O(\alpha\beta\mu)$ time.
When $\mathcal{R}$ has ply $\Delta$, Theorem 2.11 gives a $(3\sqrt{2},12\Delta)$-smooth quarter-hull-supersequence $\Xi$ of size $O(\Delta^{2}n)$. Thus we immediately have the following corollary of the theorems above.
@@
If $\mathcal{R}$ has ply $\Delta$, then for the $(3\sqrt{2},12\Delta)$-smooth quarter-hull-supersequence of Theorem 2.11 with its stable disks marked, then the quarter hull of $P$ can be computed in $O(\Delta^{3}m)$ time, where $m$ is the number of unstable disks in $\mathcal{R}$.
For ease of exposition, in Section 5 we first prove Theorems 2.10 and 2.11 for the case $\Delta=1$; i.e., disjoint unit disks.
In Section 3 we solve the problem of preprocessing a set of unit intervals of bounded ply in one dimension, which is needed to prove Theorems 2.10 and 2.11 in full generality for disks of bounded ply (i.e. $\Delta>1$) in Section 6.
In Section 7 we prove Theorems 2.12 and 2.13, which are decoupled from the preprocessing phase and hold for any set of disks.
2.5 Limitations & open problems
Restrictions on the supersequences.
Problem 2.15.
Note that, although we require handling duplicates in our application, Problem 2.15 is also open even when all points are unique. We believe this question is of independent interest.
Model of computation.
We assume we are working in the real RAM model of computation, which is consistent with earlier work in our framework. (For a formal discussion of the Real RAM, refer to [15].) In order to achieve linear reconstruction time, we require our supersequences to be smooth, however, in Section 4 we do show that reconstruction in $\mathbb{R}^{1}$ can be done in linear time, without this smooth restriction, in the word RAM model, where universal hashing is allowed.
General position.
For ease of exposition, throughout we assume general position of the disk centers in the preprocessing phase and general position of the true points in the reconstruction phase; in both cases for us general position means no three points (or centers) are collinear. However, because we actively duplicate regions and points in our sequences, we will not be able to assume full general position in all subroutines (e.g. the discussion in Section 4 hinges critically on the potential existence of duplicates). Our general position assumptions may be completely lifted by using symbolic perturbation [14].
3 Preprocessing Intervals for Sorting in One Dimension
In this section we consider the problem of preprocessing unit intervals of bounded ply in $\mathbb{R}^{1}$ such that the sorted order of a corresponding set of points (values) can be recovered quickly. This will be later used as a subroutine in Section 6. We remark that the problem of optimally preprocessing intervals for sorting was considered in prior work (see for example [39] and reference therein). Here though we require that our output structure is a supersequence, as this is what will be required to use it as a subroutine for our results in $\mathbb{R}^{2}$.
Definition 3.1.
Let $\mathcal{R}$ be a set of $n$ intervals. A sorting-supersequence of $\mathcal{R}$ is a sequence $\Xi$ of (possibly reoccurring) intervals of $\mathcal{R}$ such that, for any point set $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$, the sorted order of $P$ is a subsequence of $\Xi$.
While this definition works for any set of intervals, we will focus on unit intervals of ply $\Delta$ in this work. Furthermore, we will consider an additional restriction on the sequences, and we will show that such a restricted sequence can both be computed on a set of unit intervals of ply $\Delta$, and used to recover the sorted order; whether this restriction is necessary is left as an open problem (and may depend on the model of computation; refer to Section 4).
Definition 3.2.
A sequence of values $X=x_{1},\ldots,x_{n}$ is $(\alpha,\beta)$-smooth if:
  • for any two elements $x_{i},x_{j}\in X$ such that $i<j$ but $x_{i}>x_{j}$, we have $x_{i}-x_{j}\leq\alpha$, and (distance property)
  • for any interval $J\subset\mathbb{R}$ with $|J|>1$, there are $\leq\beta|J|$ distinct elements from $X$ in $J$. (packing property)
We will also say a sorting-supersequence $\Xi$ is $(\alpha,\beta)$-smooth if the corresponding sequence of points (values) is $(\alpha,\beta)$-smooth for any $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$.
The goal of this section is to prove the following theorems.
Theorem 3.3.
Let $\mathcal{R}$ be a set of $n$ unit intervals of ply $\Delta$. There exists a $(3,2\Delta)$-smooth sorting-supersequence $\Xi$ of $\mathcal{R}$ of size $O(n\Delta)$. Furthermore, such a sequence can be computed in $O(n(\Delta+\log n))$ time.
Theorem 3.4.
Let $\mathcal{R}$ be a set of $n$ unit intervals. Given an $(\alpha,\beta)$-smooth sorting-supersequence $\Xi$ of $\mathcal{R}$, and given a point set $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$, the sorted order of $P$ can be computed in $O(|\Xi|\log(\alpha\beta))$ time.
When $\mathcal{R}$ has ply $\Delta$, Theorem 3.3 gives a $(3,2\Delta)$-smooth sorting-supersequence $\Xi$ of size $O(n\Delta)$. Thus we immediately have the following corollary of Theorem 3.4.
Corollary 3.5.
If $\mathcal{R}$ has ply $\Delta$, then for the $(3,2\Delta)$-smooth sorting-supersequence of Theorem 3.3, the sorted order of $P$ can be computed in $O(n\Delta\log(\Delta))$ time.

### v2

@@ 2.1 Computation of the Disk Classification
Suppose a guaranteed disk $D$ appears twice on the strip hull. Then there must be two other disks $A$ and $B$ such that there are four points $a\in A$, $b\in B$, and $d_{1},d_{2}\in D$ on the convex hull of $\mathcal{R}$ belonging to the disks in the order $a,d_{1},b,d_{2}$. Let $m$ be the intersection point of $ab$ with $d_{1}d_{2}$. If we place the point of $A$ at $a$, the point of $B$ at $b$, and the point of $D$ at $m$, $D$ does not contribute a point to the convex hull, a contradiction.
Lemma 2.6.
Let $D$ be a guaranteed boundary disk. Then $D$ is stable if and only if both of its adjacent strips on the strip hull are empty, and both neighbors of $D$ on the strip hull are guaranteed.
Proof 2.7.
Let $D$ be a guaranteed boundary disks. By Observation 2.1, $D=B_{i}$ for a unique $i$.
We will first argue the $\Leftarrow$ direction. Suppose $B_{i-1}$ and $B_{i+1}$ are both guaranteed, and the strips $s_{i-1}$ and $s_{i}$ are empty (that is, do not intersect any other disks). By definition of guaranteed, the three disks $B_{i-1},B_{i},B_{i+1}$ will contribute three points $p_{i-1},p_{i},p_{i+1}$ to the convex hull. Furthermore, there can be no other point on the hull between $p_{i-1}$ and $p_{i}$ or between $p_{i}$ and $p_{i+1}$, since such a point would need to lie in $s_{i-1}$ or $s_{i}$, which is impossible. That is, $p_{i-1}$, $p_{i}$ and $p_{i+1}$ appear consecutively on the hull for all possible realizations, which implies that $p_{i}$ is indeed stable.
Next, we will argue the $\Rightarrow$ direction. Suppose that $B_{i}$ is stable. Suppose further for contradiction that $B_{i+1}$ is not guaranteed (the argument for $B_{i-1}$ is symmetric). Then there are two realizations $P$ and $P^{\prime}$ such that $p_{i+1}$ is a hull vertex in $P$, but not in $P^{\prime}$. Since disks are continuous connected regions, we may continuously move the points from $P$ to $P^{\prime}$, and we must reach some intermediate realization $P^{\prime\prime}$ in which $p_{i+1}$ is exactly on the convex hull boundary. Now, moving $p_{i}$ inside a small neighbourhood will make $p_{i+1}$ be on or not on the convex hull, contradicting that $B_{i}$ is stable. Similarly, assume for contradiction that $s_{i}$ is not empty (the argument for $s_{i-1}$ is symmetric). Then there exists a realization $P$ with a point $p_{j}$ inside $s_{i}$; draw a line through $p_{j}$ that intersects both $B_{i}$ and $B_{i+1}$ and place $p_{i}$ and $p_{i+1}$ on this line. We may assume $p_{i}$, $p_{j}$, and $p_{i+1}$ are all on the convex hull (if not, there must be another point in $s_{i}$ and we can move $\ell$ parallel to itself until it intersects this point; repeat if necessary). But then again, moving $p_{i}$ inside an epsilon neighbourhood will cause $p_{j}$ to be on the hull or not, contradicting that $B_{i}$ is stable.
Now we are ready to prove the following.
@@
Proof 2.9.
First, the boundary disks can be easily detected in $O(n\log n)$ time by computing the convex hull of the disks [12]. Second, Observation 2 allows us to classify the disks into guaranteed, potential, and impossible disks using a simple $O(n\log n)$ algorithm [36]. By Observation 2 all impossible interior disks are stable. It remains to distinguish the stable from the unstable guaranteed boundary disks.
By Lemma 2.6, in order for any guaranteed boundary disk $B_{i}$ to be stable, its adjacent strips must be empty and its neighbors on the strip hull must be guaranteed. As for the running time for finding the stable guaranteed boundary disk, as discussed above we can already determine all boundary disks, and for each such disks whether it and its neighbors are guaranteed, in $O(n\log n)$ time. Thus we only determine whether the adajacent strips are empty, though this can be done for all stips in $O(n\log n)$ time using Lemma 4.21.444Technically that lemma assumes unit disks, though the same arguments hold for non-unit disks. Namely, the strips are still pseudodisks as the spines form a plane geometric graph.
3 The Preprocessing Framework: Definitions, Preliminaries, and Results
In this section we formally define the problem and state our result for preprocessing unit disks of bounded ply in $\mathbb{R}^{2}$ for reconstructing the convex hull. While our focus is on unit disks in this paper, many of the definitions and observations in this section hold more generally for any set of disks. Before we can state our results, we need to first discuss a notion of quarter hull to decompose the problem into smaller parts, and a notion of smoothness.
@@ 3.1 Quarter Hulls
Note that the convex hull is cyclic in nature: it has no well-defined start or end point, and in the presence of uncertainty, it may be difficult even to designate an arbitrary disk as the start point, as there may not be any guaranteed disks.
To make life easier, we opt to focus on only one quarter of the convex hull: the part from the rightmost point on the hull to the topmost point on the hull (i.e. the positive quadrant). Indeed, the idea of splitting the computation of the convex hull into separate pieces is well established [3], and choosing quarters will have the added benefit that within such a quarter, no vertex can have an internal angle smaller than $90^{\circ}$, which will aid our geometric arguments later.
@@
In the remainder of this work, we also adapt the notation established in Section 2 to the notion of quarter hulls. That is, a disk will be called guaranteed or impossible when it is guaranteed or impossible to contribute to the quarter hull, and it will be stable when it does not influence the combinatorial structure of the quarter hull. Refer to Figure 6.
Figure 6: A set of disjoint unit disks and their types with respect to the top right quarter hull.
Stable disks for the convex hull are not necessarily stable for the quarter hull, but not vice versa. However, observe that the total number of unstable disks for the entire convex hull is at most a constant number more than the sum of the number of unstable disks in the four quarter hull problems; hence, dividing the problem into quarters does not jeopardize our aim to achieve sublinear reconstruction times.
{observation}
Let $\mathcal{R}$ be a set of disks in the plane. Let $m$ be the total number of unstable disks in $\mathcal{R}$ for the convex hull. Let $m_{1},m_{2},m_{3},m_{4}$ be the numbers of unstable disks in $\mathcal{R}$ for the four quarter hulls. Then $m<m_{1}+m_{2}+m_{3}+m_{4}<4m$.
@@ 3.2 Supersequences
Let $\mathcal{R}$ be a set of $n$ regions in the plane. A quarter-hull-supersequence of $\mathcal{R}$ is a sequence $\Xi$ of (possibly reoccurring) regions of $\mathcal{R}$ such that, for any point set $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$, the quarter hull of $P$ is a subsequence of $\Xi$.
We will focus on the case where $\mathcal{R}$ is a set of unit disks of bounded ply $\Delta$ in this work, where the ply of a set of regions is the maximum number of regions that any point in the plane is contained in. Replacing the disks in a quarter-hull-supersequence by their points, we now obtain a sequence of points that contains all points of the quarter hull of $P$ in the correct order. Since it is not clear if the convex hull of any such sequence can be computed in linear time (Open Problem 1.2), we additionally introduce smooth sequences.
For two points $p,q\in\mathbb{R}^{2}$, let $\nwarrow(p,q)$ be the signed distance from $p$ to $q$ projected on a line with slope 1-1; that is,
@@
Essentially, a smooth supersequence may have points that are out of order, but the extent in which they can be out of order is limited by the parameters $\alpha$ and $\beta$, which makes the reconstruction problem significantly easier.
@@
3.3 Problem Statement & Results
We now state our general problem of preprocessing disks in the plane for the convex hull, and then afterwards state our results.
Problem 3.3.
Let $\mathcal{R}$ be a set of $n$ disks. Can we construct a supersequence $\Xi$ of $\mathcal{R}$ such that given any $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$ we can construct $\mathrm{CH}(P)$ in $O(n)$ time by using $\Xi$?
While the above problem is stated for a general set of disks, some restrictions will be required. For instance, if we allow $n$ copies of the same disk, then constructing such a sequence $\Xi$ is not possible.
We now state our main results for preprocessing unit disks in the plane of bounded ply for the convex hull. Our first two theorems imply that the classical result of $O(n\log n)$ preprocessing for $O(n)$ reconstruction (as in Figure 1) can indeed be replicated for unit disks of bounded ply using supersequences.
@@ 3.3 Problem Statement & Results
Let $\mathcal{R}$ be a set of $n$ disks in the plane of constant ply and a constant range of sizes. There exists a $(O(1),O(1))$-smooth quarter-hull-supersequence $\Xi$ of $\mathcal{R}$ of size $O(n)$, and it can be computed in $O(n\log n)$ time.
Note the following theorems do not require the ply or size range to be constant.
@@
Let $\mathcal{R}$ be a set of $n$ disks. Given an $(\alpha,\beta)$-smooth quarter-hull-supersequence $\Xi$ of $\mathcal{R}$, and given a point set $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$, the quarter hull of $P$ can be computed in $O(\alpha\beta|\Xi|)$ time.
The next theorem implies that we can also obtain sublinear reconstruction using this method, which is a benefit of our approach of directly studying the convex hull, as opposed to the convex hull being product of some other geometric structure of study.
@@
Let $\mathcal{R}$ be a set of $n$ disks. We are given an $(\alpha,\beta)$-smooth quarter-hull-supersequence $\Xi$ of $\mathcal{R}$ in which all its stable disks have been marked, and let $\mu$ denote the number of unstable disks in $\Xi$. Then given a point set $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$, the quarter hull of $P$ can be computed in $O(\alpha\beta\mu)$ time.
Note for the above theorem, that by nature of sublinear reconstruction we cannot spend time proportional to the output size, therefore the algorithm will output a pointer to the head of a linked list that contains all the elements of the quarter hull in the correct order. See Section 5.2 for details.
When $\mathcal{R}$ has constant ply and constant size range, Theorem 3.4, Lemma 2.8, Theorem 3.6, and Observation 3.1 together imply the following corollary.
@@
Let $\mathcal{R}$ be a set of $n$ disks in $\mathbb{R}^{2}$ with constant ply and constant size range. We can preprocess $\mathcal{R}$ in $O(n\log n)$ time into a supersequence with its stable disks marked, from which the convex hull of $P$ can be computed in $O(m)$ time, where $m$ is the number of unstable disks in $\mathcal{R}$.
4 The Preprocessing Phase